# Riccardo Giordana Pozzi

### v1

@@ 3 Universal sectors
The purpose of this section is to identify the operators that can be exchanged in the superdisplacement four-point functions, up to NLO. Building on the considerations of subsection 2.2, our goal is to determine whether these correlators exhibit universal behavior. By universal we mean that, independently of the details of the ambient theory, such as the amount of supersymmetry, the four-point function of operators with a given conformal dimension always takes the same functional form, modulo eventual tensor structures.
As we will see, universality ultimately traces back to the constraint that the leading order must correspond to a GFF theory and that the spectrum is restricted to the displacement supermultiplet and its composite operators. As discussed in subsection 2.2, one can exploit the leading order data of the GFF to identify the $\epsilon_{\mathcal{T}}$-order of the exchanged operators at NLO. This allows to reduce the possible operators that can be exchanged in a given four-point function. Indeed, in principle, all operators compatible with symmetry can appear. However, if one can demonstrate that only operators built from the external fields are actually exchanged in the four-point, while all other operators that could in principle appear are excluded, then the correlator becomes independent of the other supermultiplet components, thereby exhibiting universality.
@@
The first requirement implies that in all the case we will consider the leading order as completely defined by taking Wick contractions of the operator insertions. Moreover, being a $1d$ dCFT, the ${\mathfrak{so}}(1,2)_{\text{conf}}$ subalgebra will always be present regardless from the case in hand, hence one can always express the four-point functions as a sum of $\mathfrak{s}\mathfrak{l}(2,\mathbb{R})$ blocks [^4] This is true even in the case in which one considers superblocks, as they can always be decomposed in a finite sum of conformal blocks.. The second condition implies that we exclude all theories having additional contributions coming from other elementary fields, e.g. tilt supermultiplets, that could in principle lead to non trivial interactions with the displacement multiplet. The last requirement is a simplifying assumption. While it may appear specific, we will see that many theories of interest naturally satisfy this property. The implications of relaxing these assumptions will be discussed at a later stage.
@@ 3.1 Generalized Free Field
We therefore begin with the analysis by studying the leading-order generalized free field theory, where correlation functions can be computed via Wick contractions of the $n$-point insertions. The lack of interactions makes this framework particularly restrictive. In fact, the only non-vanishing three-point functions involving two fundamental fields are those in which the third operator is built from the fundamental field itself. In particular we are interested in
@@
We use the notation $\phi]^{L}_{M}$ for composite operators built from the fundamental operator $\phi$, of length $L$ and with a total of $M$ derivatives. We will not be interested in the particular expression of these operator [^5] Even at the level of a GFF computation, generic operators such as $\mathbb{D}\partial^{n}_{t}\mathbb{D}$ are not primaries in one dimension, and one must construct appropriate linear combinations to obtain them. See [18] for the explicit realizations., but rather on the length of these. Working in such a basis, which in general is not the same in which the dilatation operator is diagonal, allows in general to have a better control on correlation functions of operator having a well-defined length, see [19] for a complete discussion.
@@
we conclude that in the displacement four-point function at leading order, (2.10), only multiparticle operators of length $L=2$ built out of the displacement can be exchanged. All other exchange do not appear at leading order. A similar analysis holds also for all the other four-point functions in the multiplet.
Therefore, one can conclude that the sectors are actually decoupled and so the expression for $f_{\mathbb{P}^{4}}^{(0)}(z)$, $f_{\mathbb{O}^{4}}^{(0)}(z)$ and $f_{\mathbb{D}^{4}}^{(0)}(z)$ are hence universal [^6] We stress that these statements are true modulo global symmetries. These, indeed, can introduce tensorial features dressing the functions with indexes. For example, if the displacement is charged under a global symmetry, e.g. $\mathbb{D}\rightarrow\mathbb{D}^{a}$ then $f_{\mathbb{D}^{4}}(z)\rightarrow f_{\mathbb{D}^{4}}^{a_{1}a_{2}a_{3}a_{4}}(z)$. However, one can always decompose to relate two different four-point function, see subsection 4.2. Clearly, this is a well known fact of GFF. In the following, we will look up to which extent this feature is preserved at next-to-leading order.
@@ 3.2 Next-to-leading order
We can now proceed with a similar analysis by considering the first non-trivial contributions arising from interactions. Just for reference, we will focus in the following in the four-point of the displacement operator, which, as discussed, is always present in the multiplet as it is related to the breaking of orthogonal translations. However, the same analysis can be performed for each component of a given supermultiplet.
@@
where the dots represent other exchange of which we are now not interested in. When taking into account the expansion in $\epsilon_{\mathcal{T}}$ of both sides, as discussed in section 2, one gets and expression like (2.14). Given that $\lambda_{\mathbb{D}\mathbb{D}[\mathbb{O}]_{M}^{L=2}}^{(0)}=0$, the relevant contribution associated to these exchanges at NLO is given by
@@
Exploiting the discussion presented in subsection 2.2 and knowing that $\lambda_{\mathbb{O}\mathbb{O}[\mathbb{O}]_{M}^{L=2}}^{(0)}\sim{\mathcal{O}}(1)$, consistency with the $\epsilon_{\mathcal{T}}$-order implies that
@@
Before moving on to cases where some of the assumptions are relaxed, it is worth commenting on the mixing problem. A key advantage of the analysis presented in subsection 2.2 and in the present discussion is that the identification of universal sectors provides direct control over mixing: it allows us to exclude not only mixing between different components of the same multiplet, but also the exchange of higher-length operators within each sector. As a result, the exchanged operators up to NLO are restricted to multiparticle operators with length $L=2$.
@@ 3.2.1 Relaxing the third assumption
Much of the discussion so far has relied on the identification of non-vanishing GFF three-point functions. Combined with the vanishing of the three-point function involving two displacement operators and a third one, this allowed us to constrain the minimal order at which the latter interaction can appear. While many of these constraints followed from consistency, the condition in (3.4) was instead imposed as an assumption.
@@
In the CFT language, this translates into the fact that a given four-point function also exchanges other fundamental operators belonging to the supermultiplet. Being these theory dependent, the resulting four-point functions will in general differ from one theory to another. Notice, however, that the discussion regarding multiparticle operators remains the same as in the previous case. It then follows that the only source of mixing between different sectors is precisely the one described above.
@@
Consider, for example, that one finds $\lambda_{\mathbb{D}\mathbb{D}\mathbb{O}}=0$ and $\lambda_{\mathbb{D}\mathbb{D}\mathbb{D}}\neq 0$ as arising from symmetry discussion. In that case, the displacement operator four-point function would again receive no contributions from other sectors. Nevertheless, this situation would still differ from the previous one, precisely because of the additional allowed interaction. In these case, the expansion (2.8) will factorize as
@@
Although the discussion has been framed around the displacement operator, it applies more generally to any other choice of defect operator in the multiplet. This will be, for example, the case of the $1/2$ BSP Wilson line in ${\mathcal{N}}=2$ gauge theories, [20], discussed in subsection 4.3 where is the superprimary that disentangle from the rest of the multiplet.
@@ 3.2.2 The $\mathbb{Z}_{2}$-odd sectors
We aim to better understand the operator content exchanged at NLO when the $\mathbb{Z}_{2}$ symmetry is broken. As in the $\mathbb{Z}_{2}$-preserving case, where length $L=2$ operators are exchanged at leading and subleading order, we may expect additional contributions from the $\mathbb{Z}_{2}$-odd sector. In particular, the fundamental operators associated with non-vanishing three-point functions are naturally expected to appear, but higher-length operators may also be exchanged as a consequence. To keep the discussion clearer, we restrict to displacement correlators and assume that they receive no contributions from the other sector. In this setup, we therefore allow for a non-vanishing three-point function
@@
Unlike the $\mathbb{Z}_{2}$-even case, where contributions can be analyzed via Wick contractions in the free theory, no analogous procedure exists for the $\mathbb{Z}_{2}$-odd sector. Indeed, the $\mathbb{Z}_{2}$-odd $n$-point functions vanish in the GFF and become non-vanishing only when interactions are included. Consequently, one cannot directly exploit the GFF analysis to study these three-point functions. The operators appearing in this context are therefore inherently interacting and cannot, in general, be regarded as simple products of fundamental fields, unlike in the GFF case.
@@
where we subtract all the operators that can appear in the nested OPEs so that they compensate all the divergences that can appear from the limits in the point-splitting. The advantage of exploiting these expression is that it allows us to easily relate the OPE coefficients of $\mathbb{D}]^{L}_{n}(t)$ to those of the OPE of its components and extend (3.9) also away from the GFF. Moreover, one does not actually need the entire OPE, as higher-dimensional operators will just contributes with regular term. In order to make sense of the regularization in (3.10), we focus, for clarity, on the simplest multiparticle operator $\mathbb{D}]^{L=3}_{0}(t)$, noting that the generalization to arbitrary $L$ follows similar steps. In this regard, let us first consider the three-point function
@@
where $\sigma^{ijk}$ account for all the permutations of the OPE couples and $\mathbb{T}$ are the $\mathbb{Z}_{2}$-odd operators appearing in $\mathbb{D}\times\mathbb{D}$. The $k_{\mathbb{D}\mathbb{D}\mathbb{T}}$ coefficients give the contribution that subtract the divergences and $(\mathbb{n}_{\mathbb{D}}C_{\mathcal{T}})$ is the two point normalization that comes from the identity exchange. From this expression we can see that only operators with $\Delta_{\mathbb{T}}<3\Delta_{\mathbb{D}}$ give contributions that are divergent and compensate the one that arise from the point splitting. We consider now the three-point function
@@
Under the assumption that this sector is decoupled from the others, we can see that the only contributions in $\mathbb{D}\times\mathbb{D}$ that are $\mathbb{Z}_{2}$-odd are those coming from the conformal family of $\mathbb{D}$. By taking the contractions and comparing with (3.11) we identify
@@
By similar arguments one can study all the higher-length operators. However, these will require a higher-number of expansions thus leading to higher powers of $\lambda^{(1)}_{\mathbb{D}\mathbb{D}\mathbb{D}}$ and hence negligible at NLO. Therefore only multiparticle operators that can actually be exchanged in the four-point function up to NLO are the $L=3$ ones. Moreover, since we show that their three-point function is controlled by $\lambda_{\mathbb{D}\mathbb{D}\mathbb{D}}^{(1)}$, by setting this to zero remove all the $\mathbb{Z}_{2}$-odd exchange to the four-point function. This is why we can consider the NLO function $f^{(1)}(z)$ as factorized in $\mathbb{Z}_{2}$-even and $\mathbb{Z}_{2}$-odd part (3.7), with the $\mathbb{Z}_{2}$-odd contribution controlled by $\lambda_{\mathbb{D}\mathbb{D}\mathbb{D}}$.
As usual, here we focus on the $\mathbb{D}$-sector, however, this discussion can be generalizes to all the other possibilities and also for the mixed composite multiparticle operators. For example one may expect that if we have mixing with another sector, e.g. $\mathbb{O}\in\mathbb{D}\times\mathbb{D}$, then
@@ 3.2.3 Relaxing the second assumption
In some cases, the theory contains additional fundamental operators, which we denote by ${\mathcal{T}}_{i}$ for some $i\in\mathbb{N}$. One must then consider all possibilities for whether $\mathcal{D}$} and ${\mathcal{T}}_{i}$ appear in the OPEs ${\mathcal{D}}\times{\mathcal{D}}$ and ${\mathcal{T}}_{i}\times{\mathcal{T}}_{j}$. By repeating the analysis outlined previously and using arguments similar to those in subsection 2.2, one can determine the minimal order at which exchanged operators appear and identify whether a given four-point function exhibits universal features.
@@ 3.2.4 Relaxing the first assumption
For all the cases discussed here and in the following, relaxing the first assumption will not be of particular utility. However, we discuss this possibility, leaving an in depth study for future works.
First of all, the first thing one could do is to consider still a GFF theory at leading order, but instead of a $1d$ dCFT, one could study a higher dimensional defect. As Wilson lines provides a realization for line defect, surface defects finds many interesting applications. A known example is the $1/2$ BPS surface in $6d$ $(2,0)$ theory at large $N$ [30, 33]. The main philosophy of the discussion presented here may be extended to discuss also these cases. Moreover, the knowledge of this theory up to NLO order, provides a stating point to study other surface defects by the identification of the universal sector.
Another way to relax the first assumptions is to consider the leading order not to be a GFF. In this case, the whole analysis discussed in this section will not apply. However, by studying the new OPE data of the leading order, one may still consider the discussion presented in subsection 2.2 to study and constraint the NLO order OPE that are allowed by consistency with the strong coupling expansion.

### v2

@@ 3 Universal sectors
The purpose of this section is to identify the operators that can be exchanged in the superdisplacement four-point functions, up to NLO. Building on the considerations of Sec. 2.2, our goal is to determine whether these correlators exhibit universal behavior. By universal we mean that, independently of the details of the ambient theory, such as the amount of supersymmetry, the four-point function of operators with a given conformal dimension always takes the same functional form, modulo eventual tensor structures.
As we will see, universality ultimately traces back to the constraint that the leading order must correspond to a GFF theory and that the spectrum is restricted to the displacement supermultiplet and its composite operators. As discussed in Sec. 2.2, one can exploit the leading-order data of the GFF to identify the $\epsilon_{\mathcal{T}}$ order of the exchanged operators at NLO. This allows one to reduce the possible operators that can be exchanged in a given four-point function. Indeed, in principle, all operators compatible with symmetry can appear. However, if one can demonstrate that only operators built from the external fields are actually exchanged in the four-point, while all other operators that could in principle appear are excluded, then the correlator becomes independent of the other supermultiplet components, thereby exhibiting universality.
@@
The first requirement implies that in all the case we will consider the leading order as completely defined by taking Wick contractions of the operator insertions. Moreover, being a $1d$ dCFT, the ${\mathfrak{so}}(1,2)_{\text{conf}}$ subalgebra will always be present regardless from the case in hand, hence one can always express the four-point functions as a sum of $\mathfrak{s}\mathfrak{l}(2,\mathbb{R})$ blocks. [^5] This is true even in the case in which one considers superblocks, as they can always be decomposed in a finite sum of conformal blocks. The second condition implies that we exclude all theories having additional contributions coming from other elementary fields, e.g. tilt supermultiplets, that could in principle lead to nontrivial interactions with the displacement multiplet. The last requirement is a simplifying assumption. While it may appear specific, we will see that many theories of interest naturally satisfy this property. The implications of relaxing these assumptions will be discussed at a later stage.
@@ 3.1 Generalized Free Field
We therefore begin with the analysis by studying the leading-order generalized free field theory, where correlation functions can be computed via Wick contractions of the $n$-point insertions. The lack of interactions makes this framework particularly restrictive. In fact, the only nonvanishing three-point functions involving two fundamental fields are those in which the third operator is built from the fundamental field itself. In particular we are interested in
@@
We use the notation $\phi]^{L}_{M}$ for composite operators built from the fundamental operator $\phi$, of length $L$ and with a total of $M$ derivatives. We will not be interested in the particular expression of these operators [^6] Even at the level of a GFF computation, generic operators such as $\mathbb{D}\partial^{n}_{t}\mathbb{D}$ are not primaries in one dimension, and one must construct appropriate linear combinations to obtain them. See [34] for the explicit realizations., but rather on the length of these. Working in such a basis, which in general is not the same in which the dilatation operator is diagonal, allows in general to have a better control on correlation functions of operator having a well-defined length; see [35] for a complete discussion.
@@
we conclude that in the displacement four-point function at leading order, (2.11), only multiparticle operators of length $L=2$ built out of the displacement can be exchanged. All other exchanges do not appear at leading order. A similar analysis holds also for all the other four-point functions in the multiplet.
Therefore, one can conclude that the sectors are actually decoupled and so the expressions for $f_{\mathbb{P}^{4}}^{(0)}(z)$, $f_{\mathbb{O}^{4}}^{(0)}(z)$ and $f_{\mathbb{D}^{4}}^{(0)}(z)$ are hence universal. [^7] We stress that these statements are true modulo global symmetries. These, indeed, can introduce tensorial features dressing the functions with indexes. For example, if the displacement is charged under a global symmetry, e.g. $\mathbb{D}\rightarrow\mathbb{D}^{a}$ then $f_{\mathbb{D}^{4}}(z)\rightarrow f_{\mathbb{D}^{4}}^{a_{1}a_{2}a_{3}a_{4}}(z)$. However, one can always decompose to relate two different four-point functions; see Sec. 4.2 Clearly, this is a well-known fact of GFF. In the following, we will look up to which extent this feature is preserved at next-to-leading order.
@@ 3.2 Next-to-leading order
We can now proceed with a similar analysis by considering the first nontrivial contributions arising from interactions. Just for reference, we will focus in the following on the four-point of the displacement operator, which, as discussed, is always present in the multiplet as it is related to the breaking of orthogonal translations. However, the same analysis can be performed for each component of a given supermultiplet.
@@
where the dots represent other exchange of which we are now not interested in. When taking into account the expansion in $\epsilon_{\mathcal{T}}$ of both sides, as discussed in Sec. 2, one gets and expression like (2.16). Given that $\lambda_{\mathbb{D}\mathbb{D}[\mathbb{O}]_{M}^{L=2}}^{(0)}=0$, the relevant contribution associated to these exchanges at NLO is given by
@@
Exploiting the discussion presented in Sec. 2.2 and knowing that $\lambda_{\mathbb{O}\mathbb{O}[\mathbb{O}]_{M}^{L=2}}^{(0)}\sim{\mathcal{O}}(1)$, consistency with the $\epsilon_{\mathcal{T}}$ order implies that
@@
Before moving on to cases where some of the assumptions are relaxed, it is worth commenting on the mixing problem. A key advantage of the analysis presented in Sec. 2.2 and in the present discussion is that the identification of universal sectors provides direct control over mixing: it allows us to exclude not only mixing between different components of the same multiplet, but also the exchange of higher-length operators within each sector. As a result, the exchanged operators up to NLO are restricted to multiparticle operators with length $L=2$.
@@ 3.2.1 Relaxing the third assumption
Much of the discussion so far has relied on the identification of nonvanishing GFF three-point functions. Combined with the vanishing of the three-point function involving two displacement operators and a third one, this allowed us to constrain the minimal order at which the latter interaction can appear. While many of these constraints followed from consistency, the condition in (3.4) was instead imposed as an assumption.
@@
In the CFT language, this translates into the fact that a given four-point function also exchanges other fundamental operators belonging to the supermultiplet. Being that these are theory dependent, the resulting four-point functions will in general differ from one theory to another. Notice, however, that the discussion regarding multiparticle operators remains the same as in the previous case. It then follows that the only source of mixing between different sectors is precisely the one described above.
@@
Consider, for example, that one finds $\lambda_{\mathbb{D}\mathbb{D}\mathbb{O}}=0$ and $\lambda_{\mathbb{D}\mathbb{D}\mathbb{D}}\neq 0$ as arising from symmetry discussion. In that case, the displacement operator four-point function would again receive no contributions from other sectors. Nevertheless, this situation would still differ from the previous one, precisely because of the additional allowed interaction. In this case, the expansion (2.9) will factorize as
@@
Although the discussion has been framed around the displacement operator, it applies more generally to any other choice of defect operator in the multiplet. This will be, for example, the case of the $1/2$ BSP Wilson line in ${\mathcal{N}}=2$ gauge theories, [40], discussed in Sec. 4.3 where it is the superprimary that disentangles from the rest of the multiplet.
@@ 3.2.2 The $\mathbb{Z}_{2}$-odd sectors
We aim to better understand the operator content exchanged at NLO when the $\mathbb{Z}_{2}$ symmetry is broken. As in the $\mathbb{Z}_{2}$-preserving case, where length $L=2$ operators are exchanged at leading and subleading order, we may expect additional contributions from the $\mathbb{Z}_{2}$-odd sector. In particular, the fundamental operators associated with nonvanishing three-point functions are naturally expected to appear, but higher-length operators may also be exchanged as a consequence. To keep the discussion clearer, we restrict our discussion to displacement correlators and assume that they receive no contributions from the other sector. In this setup, we therefore allow for a nonvanishing three-point function
@@
Unlike the $\mathbb{Z}_{2}$-even case, where contributions can be analyzed via Wick contractions in the free theory, no analogous procedure exists for the $\mathbb{Z}_{2}$-odd sector. Indeed, the $\mathbb{Z}_{2}$-odd $n$-point functions vanish in the GFF and become nonvanishing only when interactions are included. Consequently, one cannot directly exploit the GFF analysis to study these three-point functions. The operators appearing in this context are therefore inherently interacting and cannot, in general, be regarded as simple products of fundamental fields, unlike in the GFF case.
@@
where we subtract all the operators that can appear in the nested OPEs so that they compensate all the divergences that can appear from the limits in the point-splitting. The advantage of exploiting these expression is that it allows us to easily relate the OPE coefficients of $\mathbb{D}]^{L}_{n}(t)$ to those of the OPE of its components and extend (3.9) also away from the GFF. Moreover, one does not actually need the entire OPE, as higher-dimensional operators will just contribute with regular terms. In order to make sense of the regularization in (3.10), we focus, for clarity, on the simplest multiparticle operator $\mathbb{D}]^{L=3}_{0}(t)$, noting that the generalization to arbitrary $L$ follows similar steps. In this regard, let us first consider the three-point function
@@
where $\sigma^{ijk}$ account for all the permutations of the OPE couples and $\mathbb{T}$ are the $\mathbb{Z}_{2}$-odd operators appearing in $\mathbb{D}\times\mathbb{D}$. The $k_{\mathbb{D}\mathbb{D}\mathbb{T}}$ coefficients give the contribution that subtract the divergences and $(\mathbb{n}_{\mathbb{D}}C_{\mathcal{T}})$ is the two-point normalization that comes from the identity exchange. From this expression we can see that only operators with $\Delta_{\mathbb{T}}<3\Delta_{\mathbb{D}}$ give contributions that are divergent and compensate the one that arises from the point-splitting. We consider now the three-point function
@@
Under the assumption that this sector is decoupled from the others, we can see that the only contributions in $\mathbb{D}\times\mathbb{D}$ that are $\mathbb{Z}_{2}$ odd are those coming from the conformal family of $\mathbb{D}$. By taking the contractions and comparing with (3.11) we identify
@@
By similar arguments one can study all the higher-length operators. However, these will require a higher number of expansions thus leading to higher powers of $\lambda^{(1)}_{\mathbb{D}\mathbb{D}\mathbb{D}}$ and hence negligible at NLO. Therefore only multiparticle operators that can actually be exchanged in the four-point function up to NLO are the $L=3$ ones. Moreover, since we show that their three-point function is controlled by $\lambda_{\mathbb{D}\mathbb{D}\mathbb{D}}^{(1)}$, by setting this to zero remove all the $\mathbb{Z}_{2}$-odd exchange to the four-point function. This is why we can consider the NLO function $f^{(1)}(z)$ as factorized in $\mathbb{Z}_{2}$-even and $\mathbb{Z}_{2}$-odd parts (3.7), with the $\mathbb{Z}_{2}$-odd contribution controlled by $\lambda_{\mathbb{D}\mathbb{D}\mathbb{D}}$.
As usual, here we focus on the $\mathbb{D}$-sector; however, this discussion can be generalized to all the other possibilities and also for the mixed composite multiparticle operators. For example one may expect that if we have mixing with another sector, e.g. $\mathbb{O}\in\mathbb{D}\times\mathbb{D}$, then
@@ 3.2.3 Relaxing the second assumption
In some cases, the theory contains additional fundamental operators, which we denote by ${\mathcal{T}}_{i}$ for some $i\in\mathbb{N}$. One must then consider all possibilities for whether $\mathcal{D}$} and ${\mathcal{T}}_{i}$ appear in the OPEs ${\mathcal{D}}\times{\mathcal{D}}$ and ${\mathcal{T}}_{i}\times{\mathcal{T}}_{j}$. By repeating the analysis outlined previously and using arguments similar to those in Sec. 2.2, one can determine the minimal order at which exchanged operators appear and identify whether a given four-point function exhibits universal features.
@@ 3.2.4 Relaxing the first assumption
For all the cases discussed here and in the following, relaxing the first assumption will not be of particular utility. However, we discuss this possibility, leaving an in-depth study for future works.
First of all, the first thing one could do is to consider still a GFF theory at leading order, but instead of a $1d$ dCFT, one could study a higher-dimensional defect. As Wilson lines provide a realization for line defects, surface defects find many interesting applications. A known example is the $1/2$ BPS surface in $6d$ $(2,0)$ theory at large $N$ [19, 28]. The main philosophy of the discussion presented here may be extended to discuss also these cases. Moreover, the knowledge of this theory up to NLO order provides a stating point to study other surface defects by the identification of the universal sector.
Another way to relax the first assumptions is to consider the leading order not to be a GFF. In this case, the whole analysis discussed in this section will not apply. However, by studying the new OPE data of the leading order, one may still consider the discussion presented in Sec. 2.2 to study and constrain the NLO order OPE that are allowed by consistency with the strong coupling expansion.